# STree: Speculative Tree Decoding for Hybrid State-Space Models

### v1

@@ 6 Packed 1 63
Figure 3: Left: Comparison of number of states required and number of tokens computed to get output for a full binary tree in one forward pass. STree is able to decode a packed tree sequence, while other methods need to unroll the tree into multiple sequences. Right: Runtime in milliseconds (ms) for a forward pass for different full binary trees using different algorithms.
@@
Speed 4 102.69 (1.05$\times$) 97.44 101.68 (1.12$\times$) 90.64 94.61 (1.10$\times$) 85.33 91.80 (1.16$\times$) 78.94
8 107.28 (1.09$\times$) 97.53 104.99 (1.21$\times$) 86.68 100.67 (1.31$\times$) 76.72 98.75 (1.44$\times$) 68.44
16 105.62 (1.25$\times$) 84.26 101.97 (1.49$\times$) 68.14 94.47 OOM 88.41 OOM
Memory 4 7.67 (0.91$\times$) 8.47 7.68 (0.85$\times$) 8.97 7.73 (0.83$\times$) 9.36 7.75 (0.78$\times$) 9.85
8 7.79 (0.86$\times$) 9.05 7.85 (0.80$\times$) 9.76 7.88 (0.72$\times$) 10.89 7.91 (0.64$\times$) 12.32
16 8.02 (0.75$\times$) 10.69 8.12 (0.57$\times$) 14.02 8.17 OOM 8.28 OOM
Taking one step further, we measure the generation speed of STree with packed tree input against FSS with unrolled tree input. We used a Mamba2-2.7B model as the target model and a Mamba2-130M model as the drafting model. We generate the token tree using beam search jeon2024recursive ; svirschevski2024specexec , where we perform beam search with the drafting model and keep all the tokens generated at each step of beam search, even if the beam is later discarded. This results in an N-layer tree with M tokens at each layer, where N is the number of beam search steps and M is the number of beams. We verify the target model with greedy search, where at each step, the token with the maximum conditional likelihood from the target model is compared to the corresponding child nodes in the token tree to see if we should accept the draft. The acceptance length for the two methods is the same, as the same drafted tree is used for verification. We perform this experiment on MT_Bench zhengJudgingLLMasaJudgeMTBench2023 benchmarks for generating 100 tokens. The results are presented in Tab. 1.
We can see that STree with packed input is both faster and more memory efficient than FSS with unrolled input across all tree sizes. We also notice that as the size of the tree grows, the advantage with STree becomes bigger (1.05×1.05$\times$1.05 × to 1.49×1.49$\times$1.49 × for speed and 0.91×0.91$\times$0.91 × to 0.57×0.57$\times$0.57 × for memory), which coincides with our previous finding that STree forward pass is more scalable to larger trees. Meanwhile, we note that unrolling the tree also adds overhead to the execution. Memory-wise, the increase in memory for STree is almost negligible when tree size increases, mainly due to the increase in input size. On the contrary, the repeated tokens and states required by FSS contribute to a large increase in GPU memory, making a large token tree ($N=16,M=4/5$) infeasible to compute.
@@ 5.2 End-to-end generation with hybrid models
Having verified the effectiveness of STree in performing tree decoding, we then demonstrate the potential of STree in further boosting the speed of generation for SSMs and hybrid models. We choose a Mamba2InLlama-8B [^1] Checkpoint at https://huggingface.co/JunxiongWang/Llama3.1-Mamba2-8B-distill wang2024mamba hybrid model, with 50% mix of transformers and SSMs, as the target model. As there is a lack of smaller models in the same family, we distill a 2-layer SSM from the target model using data that is used to finetune the target model. Due to hardware limitations, we use a checkpoint from 48000 steps into the distillation.
@@
Vanilla SD 67.68 (1.69$\times$) 2.04 77.18 (1.93$\times$) 2.32 78.38 (1.93$\times$) 2.34
STree 69.84 (1.74$\times$) 2.47 78.35 (1.95$\times$) 2.79 80.03 (1.98$\times$) 2.83
@@
Vanilla SD 49.24 (1.23$\times$) 1.55 52.88 (1.32$\times$) 1.65 51.80 (1.28$\times$) 1.61
STree 54.08 (1.36$\times$) 2.03 60.34 (1.50$\times$) 2.26 58.25 (1.45$\times$) 2.16
For the vanilla speculative decoding baseline, we use the 2-layer model as the drafting model to draft 1 sequence of 4 tokens every step (target input size $1\times 5$) and verify with the target model output using speculative sampling algorithm Leviathan2023Fast . For STree, we use the draft model to draft a static tree structure shown in Fig. 4(a), and use multi-step speculative sampling (MSS sampling) miao2023specinfer to verify with the target model output. Both methods use activation replay to backtrack state. We evaluate the speed of generating 1024 tokens on three different benchmarks: MT_Bench zhengJudgingLLMasaJudgeMTBench2023 , HumanEval chenEvaluatingLargeLanguage2021 , and GSM8K cobbeTrainingVerifiersSolve2021 , with two different temperatures: 0 (greedy) and 1. The results are shown in Tab. 2
We can see that across all three benchmarks used, STree is able to outperform vanilla speculative decoding. The advantage becomes more obvious when the temperature of the generation increases and the acceptance length drops as a result. We note that this performance improvement is achieved with a baseline tree generation strategy (a static tree) and a draft model still early in its distillation process. With a more advanced tree generation strategy and draft model, the acceptance length with STree will likely improve, which will translate into more generation speed improvement.
@@
We study the effect of sampling temperature on the acceptance rate and speed-up. We use the same setup as the end-to-end generation experiment with MambaInLlama model and vary the temperature used with sampling. From Fig. 4(b) and Fig. 4(c), we can see that average acceptance length drops as temperature goes up for both STree and vanilla speculative decoding, leading to a decrease in generation speed. We notice that the absolute difference between the acceptance length of STree and vanilla speculative decoding is relatively stable ($\sim 0.4$), which means as the acceptance length both drops for both STree and vanilla speculative decoding, the improvement in generation speed becomes bigger, re-affirming our results in the previous sections.
Figure 4: Left: Structure of static tree used to generate prefix token tree with the drafting model. We draft 4 steps for each iteration and 3 tokens for each layer, resulting in 13 tokens in every input sequence. Middle: Generation speed of STree and Vanilla Specultaive Decoding (SD) under different temperature. Right: Average acceptance length of STree and Vanilla Speculative Decoding (SD) under different temperature.
@@
Besides multi-step speculative sampling, we test our method with naive sampling used in miao2023specinfer ; svirschevski2024specexec . Specifically, we use the top-k tokens from the drafting distribution to build our token tree, and perform sampling with the target model. If at any token, the sampled token from the target model falls within the top-k tokens drafted, we accept that token and continue to verify the next token. We refer the readers to svirschevski2024specexec for more details. The results are shown together in Fig. 4(b) and Fig. 4(c). We can see that at low temperatures, naive sampling has a relatively high acceptance length and fast run time. As temperature increases, the acceptance length drops significantly, because the sampled token by the target model is more random and less likely to come from the top-k tokens with the drafted model. This leads to a drop in overall generation speed as well. We note that this drop in generation speed is due to the characteristic of the sampling algorithm, but not our framework. With a high acceptance length at lower temperatures, we can still gain speed improvement.
@@
We study the effect of different static tree structures on the acceptance rate and speed-up. We note that producing a wider and deeper tree also slows the drafting model and contributes to the overall runtime. The static tree configurations are shown in Fig. 5 in the Appendix. The results are presented in Tab. 3. To determine whether to use a deeper/wider tree, we need to consider not only the effect on the acceptance rate, but also the runtime of the target and drafting models. Increasing the width of the tree drastically does improve the acceptance rate (Column B vs. C), but it also slows down the models and therefore overall speed. When we increase the tree depth (Column A vs. E), we can see that the improvement in $\tau$ outweighs the runtime cost, and therefore we see an improvement. Comparing Column A and D, where the number of tokens is the same but have different connectivity, improvement can be seen in both speed and $\tau$ when a better tree structure is used. This signifies the importance of having a good tree structure in using STree. This result also agrees with our analysis on the trade-off between runtime and acceptance rate. It provides the promise that when we use a better-trained drafting model and better drafting algorithms, which would give us a better tree and token candidates, the generation speed with STree will improve.
@@
Meanwhile, hybrid models are a mix of transformer blocks and SSM blocks. As transformers have better scaling in parallel compute, we hypothesize that more transformer blocks in the model will lead to a larger improvement in generation speed by STree. We perform an ablation study using Mamba2-Llama3 models wang2024mamba with 50%, 25%, and 0% of transformer blocks, using the same drafting model we distilled. The results are shown in Tab. 4. Firstly, with autoregressive generation, models with more SSM blocks are faster, which agrees with our expectation. For generation with STree, we are still able to outperform vanilla speculative decoding using models with fewer transformer blocks. If we eliminate the effect of different acceptance lengths in the $\frac{\text{Speed-up}}{\tau}$ column, we can see that models with fewer transformers see less per acceptance length speed-up brought by STree (0.66 to 0.63), agreeing with our hypothesis.
@@
Vanilla SD 46.56 (1.17$\times$) 1.47 0.80 47.69 (1.14$\times$) 1.45 0.78 53.72 (1.22$\times$) 1.57 0.77
STree 49.04 (1.23$\times$) 1.84 0.66 48.08 (1.15$\times$) 1.77 0.64 55.32 (1.26$\times$) 2.00 0.63
@@ 6 Discussion
Our work proposes STree, an efficient algorithm to unlock the potential of speculative tree decoding for SSMs and hybrid models. We close by discussing the limitations of STree. As presented in our analysis, STree has an overhead at a short input length as compared to previous methods, and will not universally improve generation speed when applied without careful consideration. Reducing this overhead through further algorithmic innovation could lead to a better trade-off. Meanwhile, we only demonstrate improvement on 8B models, which are still considered small LLMs. scaling up our method to bigger models and more advanced speculative decoding methods may bring more societal benefits, such as energy savings for LLM inference. We leave that for the exploration of future work.

### v2

@@ 6 Packed 1 63
Figure 3: Left: Comparison of number of states required and number of tokens computed to get an output for a full binary tree in one forward pass. STree is able to decode a packed tree sequence, while other methods need to unroll the tree into multiple sequences. Right: Runtime in milliseconds (ms) for a forward pass for different full binary trees using different algorithms.
@@
[…]
Taking one step further, we measure the generation speed of STree with packed tree input against FSS with unrolled tree input. We used a Mamba2-2.7B model as the target model and a Mamba2-130M model as the drafting model. We generate the token tree using beam search jeon2024recursive ; svirschevski2024specexec , where we perform beam search with the drafting model and keep all the tokens generated at each step of beam search, even if the beam is later discarded. This results in an N-layer tree with M tokens at each layer, where N is the number of beam search steps and M is the number of beams. We verify the target model with greedy search, where at each step, the token with the maximum conditional likelihood from the target model is compared to the corresponding child nodes in the token tree to see if we should accept the draft. The acceptance length for the two methods is the same, as the same drafted tree is used for verification. We perform this experiment on the MT_Bench zhengJudgingLLMasaJudgeMTBench2023 benchmarks for generating 100 tokens. The results are presented in Tab. 2.
We can see that STree with packed input is both faster and more memory efficient than FSS with unrolled input across all tree sizes. We also notice that as the size of the tree grows, the advantage with STree becomes bigger ($1.05\times$ to $1.49\times$ for speed and $0.91\times$ to $0.57\times$ for memory), which coincides with our previous finding that STree forward pass is more scalable to larger trees. Meanwhile, we note that unrolling the tree also adds overhead to the execution. Memory-wise, the increase in memory for STree is almost negligible when tree size increases, mainly due to the increase in input size. On the contrary, the repeated tokens and states required by FSS contribute to a large increase in GPU memory, making a large token tree ($N=16,M=4/5$) infeasible to compute.
@@ 5.2 End-to-end generation with hybrid models
Having verified the effectiveness of STree in performing tree decoding, we then demonstrate the potential of STree in further boosting the speed of generation for SSMs and hybrid models. We choose a Mamba2InLlama-8B [^1] Checkpoint at https://huggingface.co/JunxiongWang/Llama3.1-Mamba2-8B-distill wang2024mamba hybrid model, with 50% mix of transformers and SSMs, as the target model. As there is a lack of smaller models in the same family, we distill a 2-layer SSM from the target model using data that is used to finetune the target model. We show the results of using a checkpoint from 48000 steps into the distillation. Results from different checkpoints are shown in the ablation study below.
@@
[…]
@@
[…]
For the vanilla speculative decoding baseline, we use the 2-layer model as the drafting model to draft 1 sequence of 4 tokens every step (target input size $1\times 5$) and verify with the target model output using speculative sampling algorithm Leviathan2023Fast. For STree, we use the draft model to draft a static tree structure shown in Fig. 4(a), and use multi-step speculative sampling (MSS sampling) miao2023specinfer to verify with the target model output. Both methods use activation replay to backtrack the state. We evaluate the speed of generating 1024 tokens on three different benchmarks: MT_Bench zhengJudgingLLMasaJudgeMTBench2023 , HumanEval chenEvaluatingLargeLanguage2021 , and GSM8K cobbeTrainingVerifiersSolve2021 , with two different temperatures: 0 (greedy) and 1. The results are shown in Tab. 3
We can see that across all three benchmarks used, STree is able to outperform vanilla speculative decoding. The advantage becomes more obvious when the temperature of the generation increases and the acceptance length drops as a result. We note that this performance improvement is achieved with a baseline tree generation strategy (a static tree) and a draft model still early in its distillation process. With a more advanced tree generation strategy and draft model, the acceptance length with STree will likely improve, which will translate into more generation speed improvement. We include the results with H100 GPUs and a beam search tree in the appendix due to space constraint.
@@
We study the effect of sampling temperature on the acceptance rate and speed-up. We use the same setup as the end-to-end generation experiment with MambaInLlama model and vary the temperature used with sampling. From Fig. 4(b) and Fig. 4(c), we can see that average acceptance length drops as temperature goes up for both STree and vanilla speculative decoding, leading to a decrease in generation speed. We notice that the absolute difference between the acceptance length of STree and vanilla speculative decoding is relatively stable ($\sim 0.4$), which means that as the acceptance length drops for both STree and vanilla speculative decoding, the improvement in generation speed becomes bigger, re-affirming our results in the previous sections.
Figure 4: Left: Structure of static tree used to generate a prefix token tree with the drafting model. We draft 4 steps for each iteration and 3 tokens for each layer, resulting in 13 tokens in every input sequence. Middle: Generation speed of STree and Vanilla Specultaive Decoding (SD) under different temperature. Right: Average acceptance length of STree and Vanilla Speculative Decoding (SD) under different temperature.
@@
Besides multi-step speculative sampling, we test our method with naive sampling used in miao2023specinfer ; svirschevski2024specexec . Specifically, we use the top-k tokens from the drafting distribution to build our token tree, and perform sampling with the target model. If at any token, the sampled token from the target model falls within the top-k tokens drafted, we accept that token and continue to verify the next token. We refer the readers to svirschevski2024specexec for more details. The results are shown together in Fig. 4(b) and Fig. 4(c). We can see that at low temperatures, naive sampling has a relatively high acceptance length and fast run time. As temperature increases, the acceptance length drops significantly, because the sampled token by the target model is less likely to come from the top-k tokens with the drafted model. This leads to a drop in overall generation speed as well. We note that this drop in generation speed is due to the characteristic of the sampling algorithm, but not our framework. With a high acceptance length at lower temperatures, we can still achieve speed improvements.
@@
We study the effect of different static tree structures on the acceptance rate and speed-up. We note that producing a wider and deeper tree also slows the drafting model and contributes to the overall runtime. The static tree configurations are shown in Fig. 5 in the Appendix. The results are presented in Tab. 4. To determine whether to use a deeper/wider tree, we need to consider not only the effect on the acceptance rate, but also the runtime of the target and drafting models. Increasing the width of the tree drastically does improve the acceptance rate (Column B vs. C), but it also slows down the models and therefore overall speed. When we increase the tree depth (Column A vs. E), we can see that the improvement in $\tau$ outweighs the runtime cost, and therefore we see an improvement. Comparing Column A and D, where the number of tokens is the same but with different connectivity, improvement can be seen in both speed and $\tau$ when a better tree structure is used. This signifies the importance of having a good tree structure in using STree. This result also agrees with our analysis on the trade-off between runtime and acceptance rate. It provides the promise that when we use a better-trained drafting model and better drafting algorithms, which would give us a better tree and token candidates, the generation speed with STree will improve.
@@
Hybrid models are a mix of transformer blocks and SSM blocks. As Transformers have better scaling in parallel compute, we hypothesize that more transformer blocks in the model will lead to a larger improvement in generation speed by STree. We perform an ablation study using Mamba2-Llama3 models wang2024mamba with 50%, 25%, and 0% of transformer blocks, using the same drafting model we distilled. The results are shown in Tab. 5. With autoregressive generation, models with more SSM blocks are faster, which agrees with our expectation. For generation with STree, we are still able to outperform vanilla speculative decoding using models with fewer transformer blocks. If we eliminate the effect of different acceptance lengths in the $\frac{\text{Speed-up}}{\tau}$ column, we can see that models with fewer transformers see less speed-up per acceptance length (0.66 to 0.63), agreeing with our hypothesis.
@@
[…]
Effect of different drafting model checkpoints
In Table 6, we provide an ablation study on our drafting model distilled to different steps. We can see that as the distillation goes on, the acceptance length for the draft model gets longer. As the acceptance length increases, the end-to-end generation speed of STree improves from being 1% slower than Vanilla SD at step 12000 to 9.67% faster at step 48000 and a further 14.8% speedup at step 264000. This is because the distribution of the drafting model is more aligned with the target model, suggesting that with better drafting model and technique that gives a better acceptance length, the end-to-end runtime of STree will become better.
Table 6: End-to-end generation speed (in Tokens/seconds) and acceptance length for Vanilla Speculative decoding and STree using drafting model trained for different steps.
12000 step 48000 step 264000 step
Speed (Vanilla SD) 44.52 (1.21x) 53.11 (1.24x) 56.49 (1.42x)
$\tau$ (Vanilla SD) 1.34 1.56 1.68
Speed (STree) 42.96 (1.19x) 58.65 (1.36x) 64.30 (1.62x)
$\tau$ (STree) 1.51 2.02 2.23
@@ 6 Discussion
Our work proposes STree, an efficient algorithm to unlock the potential of speculative tree decoding for SSMs and hybrid models. We close by discussing the limitations of STree. As presented in our analysis, STree has an overhead at a short input length as compared to previous methods, and will not universally improve generation speed when applied without careful consideration. Reducing this overhead through further algorithmic innovation could lead to a better trade-off. Meanwhile, we only demonstrate improvement on 8B models, which are still considered small LLMs. scaling up our method to bigger models and more advanced speculative decoding methods may bring more benefits, including energy savings for LLM inference. We leave that for the exploration of future work.
Acknowledgements. This work was supported by ONR award #N000142212252 and NSF 2112562 Athena AI Institute.
@@ Appendix A Static Tree configurations
Appendix B Experiments on H100 GPU
We further extend our method to H100 GPUs to demonstrate the our method is still applicable with better hardware. We apply our kernels as it is on H100 GPUs and compare it to vanilla speculative decoding.
Table 7: Generation speed (tokens/sec.) and average number of tokens accepted $\tau$ for Vanilla speculative decoding (SD) and STree with MambaInLlama-8B model with 50% mix of transformers on H100 GPU.
MT-Bench HumanEval GSM-8K
Methods Speed $\tau$ Speed $\tau$ Speed $\tau$
Temperature=0
Autoregressive 78.09 1 78.44 1 79.33 1
[…]
Temperature=1
Autoregressive 76.66 1 77.67 1 78.35 1
[…]
From the above tables, we can see that the extent of improvement from applying both Vanilla SD and STree is smaller as compared to autoregressive decoding. This is because the memory bandwidth of H100 GPU is bigger than that of RTX3090, leading to a smaller bottleneck in memory transfer. With greedy generation, STree achieves a slightly slower speed as compared to Vanilla SD. When temperature=1, STree still holds an advantage. This is due to the relative increase in acceptance length is smaller when we are doing a greedy generation and the relative overhead between the STree and Vanilla Speculative decoding increases due to the faster GPU. We believe that as the model size increases and the memory bandwidth bottleneck due to the model weight transfer becomes more significant, we will still be able to show improvement. We also note that we present a general algorithm that is not GPU specific and is not optimized for H100 GPU.
Appendix C End-to-end generation result using beam search tree
We note that STree can work with any tree generation strategy and optimizing the tree structure is orthogonal to our work. Here, we further show that the end-to-end generation results using a beam search tree for drafting, which is a basic dynamic tree structure. The expansion of the tree depends on the joint probability and will vary from sample to sample. We use greedy decoding with M = 3 and N = 4 (M, N explained in table 1). The results are shown in Table 8. We can see that with a dynamic tree generation strategy, we are able to get a better acceptance length and thus a faster end-to-end generation speed. We believe that more advance tree construction technique would bring even more benefits.
Table 8: End-to-end generation speed and acceptance length ($\tau$) for different method of drafting.
Autoregressive Vanilla SD STree w. static tree STree w. beam search tree
Speed 39.98 67.68 (1.69x) 69.84 (1.75x) 71.60 (1.79x)
$\tau$ 1 2.04 2.47 2.60